OpenStudy (anonymous):

10 less than 3 times some number is the same as the quotient of the number and 4. find the number

OpenStudy (anonymous):

x=10

OpenStudy (anonymous):

3x-10=x/4

OpenStudy (anonymous):

wrong sweetz

OpenStudy (amistre64):

im trying out firefox to see if its better; so far so good lol

OpenStudy (anonymous):

x=4

OpenStudy (anonymous):

still wrong

OpenStudy (anonymous):

40/11

OpenStudy (amistre64):

12x-40=x 11x = 40 x = 40/11 ??

OpenStudy (anonymous):

oops calc error srry

OpenStudy (anonymous):

oops calc error srry

OpenStudy (anonymous):

rush of blood to the head...

OpenStudy (anonymous):

u lisn 2 coldplay sweetz?

OpenStudy (anonymous):

nope

OpenStudy (anonymous):

gtg bye people

OpenStudy (anonymous):

bye